what is the [OH-] of a solution made by diluting 5.61 M sodium hydroxide (MM=98.076) in a final volume 500.0 mL?
1
Expert's answer
2017-05-05T06:59:57-0400

Solution:
V = 500 mL = 0.5 L
If we take 10 mL of sodium hydroxide
n(NaOH) = 5.610.01 = 0.0561 mol
C(NaOH) = n (NaOH) / final volume = 0.0561/0.5 = 0.1122 M
Sodium hydroxide is a strong electrolyte
[OH-] = C(NaOH) = 0.1122 M
If we take 100 mL of sodium hydroxide
n(NaOH) = 5.610.1 = 0.561 mol
C(NaOH) = n (NaOH) / final volume = 0.561/0.5 = 1.122 M
[OH-] = 1.122 M

Answer: [OH-] = 1.122 M